OpenStudy (anonymous):

Divide eight and two-thirds divided by two and one-third. seven twenty-sixths four and two-thirds three and five-sevenths four and one-half

OpenStudy (texaschic101):

turn the mixed numbers into improper fractions 8 2/3 divided by 2 1/3 = (26/3) / (7/3) when dividing fractions, you " flip " what you are dividing by, then multiply so...(26/3) / (7/3) = 26/3 x 3/7 = 78/21 = 3 15/21 = 3 5/7